Circular No. 7282

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999eh IN NGC 2770 G. M. Hurst, Basingstoke, England, reports the discovery by Mark

Armstrong, Rolvenden, of a supernova (mag ∼ 17.5) on an unfiltered CCD frame taken on Oct. 12.1375 UT with a 0.30-m telescope in the course of the U.K. Nova/Supernova Patrol. In response to an alert, T. Boles, Wellingborough, obtained a confirmatory image on Oct. 14, and Armstrong also secured a further image on Oct. 14.157 showing the new object at about the same brightness. Precise positions measured by Armstrong from his Oct. 12 and 14 images yield α = 9^{h}09^{m}32^{s}.67, δ = +33^{o}07′16′′.9 (equinox 2000.0), which is 20′′ west and 5′′ south of the center of NGC 2770. Armstrong’s measurement of Boles’ image yields position end figures 32^{s}.55, 16′′.3. Research by Hurst indicates that SN 1999eh is absent from the Vickers Deep Space CCD Atlas: North (1993, p. 96; limiting mag ∼ 19) and from both Palomar Sky Surveys (limiting mag ∼ 20).

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999eh, obtained by M. Calkins on Oct. 15.5 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ib supernova. The spectrum exhibits a flat continuum and broad absorption features with He I (rest 587.6, 706.5, and 492.2 nm) being most prominent. There is also evidence of weak [O I] emission, indicating that the supernova is ∼ 2 months past maximum light.

XTE J1859+226 M. L. McCollough, Universities Space Research Association and Mar-

shall Space Flight Center (MSFC); and C. A. Wilson, MSFC, report for the BATSE team: “XTE J1859+226 (IAUC 7274) was first detected with CGRO/BATSE on Oct. 9 at a flux of 0.5 Crab (9.1 × 10^{−9} erg cm^{−2} s^{−1}) and rose to 0.6 Crab (1.0 × 10^{−8} erg cm^{−2} s^{−1}) on Oct. 11, before slowly decreasing to 0.3 Crab (6.1 × 10^{−9} erg cm^{−2} s^{−1}) on Oct. 13. XTE J1859+226 has shown evidence of spectral softening in the energy band 20–100 keV, from a power-law photon index of −2.3 ± 0.1 on Oct. 9 to a photon index of −2.8 ± 0.1 at its peak (Oct. 11). XTE J1859+226 is clearly being detected up to 200 keV.”

1999 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green